# DreamMapping: High-Fidelity Text-to-3D Generation via Variational Distribution Mapping

### v1

@@ 3.2 Variational Distribution Mapping
Take rendered images as degraded cases of diffusion models. During the iterative optimization of SDS, the inferior optimization of 3D object parameters often compromises the quality of rendered images, leading to low-quality (LQ) renderings. Diffusion models face challenges in directly representing low-quality (LQ) images due to limited training in such cases. However, considering the potent capabilities of diffusion models in generating high-quality (HQ) images, it’s reasonable to assume they can approximate the HQ counterparts of rendered LQ images by treating the latter as degraded versions of the former. Specifically, given $\hat{\mathbf{x}}_{0}\sim p(x)$ as a sample of diffusion output, the rendered image $\mathbf{x}_{0}=g(\theta,c)$ can be represented by a degradation result:
@@ 3.3 Distribution Coefficient Annealing
The Jacobian term $\frac{\partial\epsilon_{\phi}(\textbf{x}_{t},t,y)-\lambda_{t}\mathcal{M}_{\psi}(\epsilon_{\phi}(\textbf{x}_{t},t,y)}{\partial\mathbf{x}_{t}}$ is also ignored following the conclusion of SDS [PJBM22].
@@ 3.4 Generation Framework
Following recent text-to-3D generation methods [PJBM22, WLW23, CCJJ23], our optimized-based framework leverages 3D Gaussian Splatting (3DGS), which is highly efficient in rendering to high-fidelity 3D creations. As Liang et al. [LYL23] have demonstrated, point cloud initialization is critical for geometry quality. We adopt Shape-E [JN23] to generate the coarse initialization with shape prior (Figure 3). Algorithm 1 shows the 3D generation process of DreamMapping.
@@ 4 Experiments
DreamFusion [PJBM22] 23.60 37.46 1.5 hours
@@ 4.4 Qualitative Results
We compared our model with current SoTA baselines using 3DGS representation (i.e., DreamGaussian [TRZ23], GaussianDreamer [YFW23], LucidDreamer [LYL23]). SD 2.1 Base was used for distillation, and all generation experiments were conducted on Nvidia A100 for fair comparison. As shown in Figure 4, we can see that DreasmGaussian suffers from low quality with blurry appearance and incomplete shape. Results of GaussianDreamer and LucidDreamer show good semantic consistency and realistic colors. However, they still generate vague textures on parts of generated models. For example, the tail light part in the case of “A supercar made out of toy bricks” is blurry, and fur details are missing in GaussianDreamer. Instead, our approach efficiently predicts an accurate variational distribution of rendered images, yielding diverse semantically meaningful 3D assets with clear geometry structure and intricate appearance details. Lastly, it is important to highlight that LucidDreamer’s generation time is typically longer than our approach due to its use of the DDIM inversion method [SME20], which requires substantially more computational resources, such as loss computations. Please refer to Figure 12 more visual results.

### v3

@@ 4.2 Variational Distribution Mapping
During the iterative optimization of SDS, the inferior optimization of 3D object parameters often compromises the quality of rendered images, leading to low-quality (LQ) renderings. Diffusion models face challenges in directly representing low-quality (LQ) images due to limited training in such cases. However, considering the potent capabilities of diffusion models in generating high-quality (HQ) images, it’s reasonable to assume they can approximate the HQ counterparts of rendered LQ images by treating the latter as degraded versions of the former. Specifically, given $\hat{\mathbf{x}}_{0}\sim p(x)$ as a sample of diffusion output, the rendered image $\mathbf{x}_{0}=g(\theta,c)$ can be represented by a degradation result:
@@ 4.3 Distribution Coefficient Annealing
The Jacobian term $\frac{\partial\epsilon_{\phi}(\textbf{x}_{t},t,y)-\lambda_{t}\mathcal{M}_{\psi}(\epsilon_{\phi}(\textbf{x}_{t},t,y)}{\partial\mathbf{x}_{t}}$ is also ignored following the conclusion of SDS [PJBM23].
@@ 4.4 Generation Framework
Following recent text-to-3D generation methods [PJBM23, WLW23, CCJJ23], our optimized-based framework leverages 3D Gaussian Splatting (3DGS), which is highly efficient in rendering to high-fidelity 3D creations. As Liang et al. [LYL23] have demonstrated, point cloud initialization is critical for geometry quality. We adopt Shape-E [JN23] to generate the coarse initialization with shape prior (Figure 3). Algorithm 1 shows the 3D generation process of DreamMapping.
@@ 5 Experiments
DreamFusion [PJBM23] 23.60 37.46 1.5 hours
@@ 5.4 Qualitative Results
We compared our model with current SoTA baselines using 3DGS representation (i.e., DreamGaussian [TRZ23], GaussianDreamer [YFW24], LucidDreamer [LYL23]). SD 2.1 Base was used for distillation, and all generation experiments were conducted on Nvidia A100 for fair comparison. As shown in Figure 4, we can see that DreasmGaussian suffers from low quality with blurry appearance and incomplete shape. Results of GaussianDreamer and LucidDreamer show good semantic consistency and realistic colors. However, they still generate vague textures on parts of generated models. For example, the tail light part in the case of “A supercar made out of toy bricks” is blurry, and fur details are missing in GaussianDreamer. Instead, our approach efficiently predicts an accurate variational distribution of rendered images, yielding diverse semantically meaningful 3D assets with clear geometry structure and intricate appearance details. Lastly, it is important to highlight that LucidDreamer’s generation time is typically longer than our approach due to its use of the DDIM inversion method [SME21], which requires substantially more computational resources, such as loss computations. Please refer to Figure 12 more visual results.